OpenStudy (anonymous):

Given the equation y − 3 = one half(x + 6) in point-slope form, identify the equation of the same line in slope-intercept form. Choose one answer. a. x − 2y = −12 b. y = one halfx c. y = one halfx + 6 d. y = one halfx + 9 Given the equation y − 3 = one half(x + 6) in point-slope form, identify the equation of the same line in standard form. Choose one answer. a. x − 2y = −12 b. y = one halfx c. y = one halfx + 6 d. y = one halfx + 9

pooja195 (pooja195):

y − 3 = 1/2(x + 6) distribute the half

OpenStudy (anonymous):

y - 3 = 1/2x + 3 then add 3 both sides y= 1/2x + 6

pooja195 (pooja195):

Thats correct :)

pooja195 (pooja195):

y − 3 = one half(x + 6) Write in standred form

OpenStudy (anonymous):

so same answer

pooja195 (pooja195):

Actually im not sure about standared form :/ i never really got that maybe @johnweldon1993 can help i hope :)

OpenStudy (anonymous):

is he on?

OpenStudy (anonymous):

anyone on?

OpenStudy (anonymous):

Can anyone help with the second question?

pooja195 (pooja195):

it should have letters so you were right

jaynator495 (jaynator495):

standard form is just numbers

jaynator495 (jaynator495):

http://prntscr.com/7fr6dt =12-17